How many gallons of a 3% salt solution must be mixed with 30 gallons of a 9% solution to obtain a 6% solution?
Answer by dnanos(83)

We need,
X gallons 3%
and
30 gallons 9%
to obtain,
30+x gallons 6%
so
because the pure salt weight before and after mixing remains equal
we have the following equation.
(3/100)*x+ (9/100)*30=(6/100)*(30+x)
multiply by 100
3x+9*30=6(30+x)
3x+270=180+6x
3x-6x=180-270
-3x=-90
3x=90
3x/3=90/3
x=30
So we need 30 gallons of 3% ...